### v1

# Testing Stochastic Block Models Based on Maximum Sampling Entry-Wise Deviations

Abstract
The stochastic block model (SBM) has been widely used to analyze network data. Various goodness-of-fit tests have been proposed to assess the adequacy of model structures. To the best of our knowledge, however, none of the existing approaches are applicable for sparse networks in which the connection probability of any two communities is of order $n^{-1}\log n$, and the number of communities is divergent. To fill this gap, we propose a novel goodness-of-fit test for the stochastic block model. The key idea is to construct statistics by sampling the maximum entry-deviations of the adjacency matrix that the negative impacts of network sparsity are alleviated by the sampling process. We demonstrate theoretically that the proposed test statistic converges to the Type-I extreme value distribution under the null hypothesis regardless of the network structure. Accordingly, it can be applied to both dense and sparse networks. In addition, we obtain the asymptotic power against alternatives. Moreover, we introduce a bootstrap-corrected test statistic to improve the finite sample performance, recommend an augmented test statistic to increase the power, and extend the proposed test to the degree-corrected SBM. Simulation studies and two empirical examples with both dense and sparse networks indicate that the proposed method performs well.
Keywords: Adjacency Matrix; Goodness-of-Fit Test; Maximum Sampling Entry-Wise Deviation; Sparse Networks; Stochastic Block Model
1 Introduction
Due to the rapid development and application of information technology, immense troves of network data have been created across various fields, such as finance and business (Katona et al. 2011 and Di Maggio et al. 2019), health care (Valente 2010; Lawyer 2015), engineering (Agarwal et al. 2013), machine learning (Musumeci et al. 2018) and social science (Tabassum et al. 2018). To make full use of network data, it is critical to understand their generation mechanisms. To this end, the stochastic block model (SBM) has been widely studied and used since the pioneering work of Holland et al. (1983); see, e.g., Snijders and Nowicki (1997), Nowicki and Snijders (2001), Bickel and Chen (2009), Rohe et al. (2011), Choi et al. (2012), Jin (2015), Zhang and Zhou (2016), and Hu et al. (2021). A good review paper on SBM can be found in Lee and Wilkinson (2019).
The SBM basically assumes that the network nodes can be decomposed into $K$ non-overlapping communities, where each node belongs to one specific community. In addition, the edge connecting any two nodes depends only on the communities in which they are located. Consider an undirected network containing $n$ nodes existing within $K$ communities. For each node $i\in\{1,\cdots,n\}$, $g(i)\in\{1,\cdots,K\}$ represents the community to which node $i$ belongs. Let $A=(a_{ij})\in\mathbb{R}^{n\times n}$ be a symmetric adjacency matrix, where $a_{ij}=1$ if two nodes $i$ and $j$ are connected and $a_{ij}=0$ otherwise. For the sake of completeness, the diagonal elements of $A$ are set to zero. We then assume that each $a_{ij}$, for $i>j$, is independently generated from the Bernoulli distribution with the probability $Q_{g(i)g(j)}$, where $0\leq Q_{g(i)g(j)}\leq 1$ is the connection probability of nodes $i$ and $j$.
When the community number $K$ is known, various approaches have been proposed to estimate the community membership vector $g=(g(1),\cdots,g(n))^{\top}$; see, e.g., the modularity-based method (Newman 2006), the profile-likelihood maximization method (Bickel and Chen 2009), the pseudo-likelihood maximization method (Amini et al. 2013), the variational method (Daudin et al. 2008), and the spectral clustering method (Rohe et al. 2011 and Jin 2015). In addition, some researchers have established the theoretical properties of the estimator of $g$ under the condition $n\min_{i,j}Q_{g(i)g(j)}\to\infty$ or even under stronger conditions; see, e.g., Choi et al. (2012), Rohe et al. (2011), Zhao et al. (2012), Sarkar and Bickel (2015), Jin (2015), Lei and Rinaldo (2015), and Zhang and Zhou (2016).
To broaden the usefulness of the SBM, it is crucial to check the model adequacy. Hence, Bickel and Sarkar (2016) used the largest eigenvalue of the centered and scaled adjacency matrix to test the Erdős-Rényi model, that is, the SBM with one community, Karwa et al. (2016) developed a finite-sample Monte Carlo method to assess the SBM. It is worth noting that the aforementioned testing methods rely on two stringent conditions: (i) the number of communities $K$ is finite, and (ii) the network is dense, such that $n\min_{i,j}Q_{g(i)g(j)}/\log n\rightarrow\infty$.
Relaxing Condition (i) to accommodate a divergent $K$, Lei (2016) proposed a goodness-of-fit test of the SBM. Specifically, he employed the largest singular value of the centered and rescaled adjacency matrix $A$, and then obtained the asymptotic null distribution under the condition $K=o(n^{1/6})$. To further relax this condition, Hu et al. (2021) developed a test statistic based on the maximum entry-wise deviations of the centered and rescaled adjacency matrix $A$, which allows the number of communities to grow with $K=o(n/\log^{2}n)$. Although the methods of Lei (2016) and Hu et al. (2021) are applicable for a divergent $K$, these two methods require the condition $\min_{i,j}Q_{g(i)g(j)}\geq c$ for a constant $10<c<$ 1 and it is expected that these two methods are inapplicable when Condition (ii) is not valid.
It is worth noting that sparse networks often appear in practical applications; see, e.g., Amini et al. (2013); Sarkar and Bickel (2015); Wang and Bickel (2017); and Jing et al. (2022). To get more insight regarding sparsity in networks, we consider two datasets given below. (A) The Sina Weibo network consists of 2,580 nodes, which was studied by Wu et al. (2022). In this network, $a_{ij}=1$ if user $i$ follows user $j$. The network density is 1.3%, calculated by the formula $\sum_{i,j}a_{ij}/(n(n-1))$. (B) The co-authorship network contains 3,607 nodes, which was analyzed by Ji and Jin (2016). The connection between any two nodes is defined as follows: $a_{ij}=1$ if and only if authors $i$ and $j$ have at least two coauthored papers. The network density is approximately 0.012%. To assess the sparsity of the above two networks, let $D_{n}$ denote the density of the network. We then calculate the ratio $nD_{n}/\log n$, resulting in a value of 4.273 for dataset (A) and 0.053 for dataset (B). Since both values of $nD_{n}/\log n$ are not large, these two datasets are expected to satisfy $D_{n}=O(n^{-1}\log n)$ and can be considered sparse.
In general, one can employ the SBM to analyze the above sparse datasets. Some researchers have developed test statistics and obtained their asymptotic null distributions in sparse networks. For example, Han et al. (2023) proposed a residual subsampling (RIRS) approach for testing and estimating $K$ with the network density being $O(n^{-1+\epsilon})$ for some arbitrarily small $\epsilon>0$. Zhang and Amini (2023) developed an adjusted chi-square statistic, which is applicable for testing SBM with an average degree no smaller than $O(\log n)$. However, the aforementioned two tests are only applicable for fixed $K$. Han et al. (2023) emphasized in their Section 2.1 that they required the mean of the adjacency matrix $A$ to be of fixed rank, which implicitly implies that $K$ is fixed. In addition, as discussed in Section 2.2 of Zhang and Amini (2023), they also required a constant value of $K$ in order to ensure the convergence of their adjusted chi-square test statistic to the standard normal distribution. Our simulation results in Section 3 confirm their limitation; both methods exhibit size distortions for large $K$. To the best of our knowledge, there is no goodness-of-fit test for evaluating the suitability of the SBM for a sparse network with a divergent $K$.
The aim of this paper is to develop a new test statistic that allows for testing the number of communities $K$ and the mapping vector $g=(g(1),\cdots,g(n))^{\top}$, without regard to the sparsity of the network or the divergence of $K$. To obtain this test, we not only allow $K$ to be divergent, but also simply require the condition, $n\min_{i,j}Q_{g(i)g(j)}\rightarrow\infty$. This is a relaxed version of the Condition (ii) mentioned earlier. The procedures of the proposed method are as follows. We first adopt the approach of Hu et al. (2021) to calculate the entry-wise deviation $\widehat{\rho}_{iv}$ for each node $i$ in community $v\in\{1,\cdots,K\}$, which is constructed based on the average of the standardized summation of $a_{ij}$ for all nodes $j\not=i$ within the community $v$. Subsequently, for any fixed community $v\in\{1,\cdots,K\}$, we randomly draw $B$ samples from $n$ nodes with replacement and then calculate $\widehat{\gamma}_{\cdot v}=\sum_{i=1}^{B}\widehat{\rho}_{iv}/\sqrt{B}$. Repeating this sampling procedure $M$ times for each of the $K$ communities, we obtain the $M$ values of $\widehat{\gamma}_{\cdot v}$, and denote them $\widehat{\gamma}_{mv}$ for $m=1,\cdots,M$. Finally, the new test statistic, $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\gamma}_{mv}|$, is constructed. By the central limit theorem, we can demonstrate that $\widehat{\gamma}_{mv}$ converges to the standard normal distribution $N(0,1)$ plus an additional term of the order $O_{p}(1/\sqrt{B\log n})$. By setting $B$ to be divergent, the additional term with the order $O_{p}(1/\sqrt{B\log n})$ becomes negligible. As a result, we expect that $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\gamma}_{mv}|$ converges to an extreme value distribution under certain conditions regardless of dense networks or sparse networks.
Although implementing our proposed test is not complicated, establishing its theoretical properties is challenging. This is because the $\widehat{\gamma}_{mv}$’s are correlated due to the replacement samples from $\widehat{\rho}_{iv}$. As a result, $\max_{m,v}|\widehat{\gamma}_{mv}|$ is no longer the maximum of a sequence of independent variables, and its asymptotic distribution is more difficult to demonstrate than that of Hu et al. (2021). In the paper, we overcome challenges and show that $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\gamma}_{mv}|$ converges to the Type-I extreme value distribution. In addition, according to the aforementioned explanations, our approach comprises Hu et al. (2021)’s method and the sampling with replacement procedure. Accordingly, our proposed method is versatile for both dense and sparse networks. In contrast, Hu et al. (2021)’s method is tailored for dense networks. Consequently, our proposed test exhibits greater robustness in practical application. Since the test statistic’s rate of convergence to the extreme distribution may be slow, we further consider a finite-sample bootstrap correction by generating bootstrap samples to improve the finite sample performance. Finally, we introduce an augmented test statistic to increase the power and extend the proposed test to the degree-corrected SBM.
The remainder of this paper is organized as follows. Section 2 introduces the proposed test and studies asymptotic null distribution, the power properties, the bootstrap-corrected method, and an augmented test statistic. The simulation studies are presented in Section 3, and they indicate that the proposed method outperforms those in Hu et al. (2021), Lei (2016), Han et al. (2023) and Zhang and Amini (2023) under sparse networks. Furthermore, our augmented bootstrap-corrected test is comparable or superior to those methods under both dense and sparse networks. Section 4 provides two real data examples with the dense and sparse networks, separately. As found in simulation studies, our method works well empirically regardless the network density. Section 5 discusses the degree-corrected SBM and provides concluding remarks. All the technical details and additional simulation results are relegated to the supplementary material.
2 Model and theoretical properties
2.1 Hypothesis testing
Consider a network of $n$ nodes with the adjacency matrix $A$. Our focus is to assess whether this network can be successfully fitted by a stochastic block model with $K_{0}$ communities and a membership vector $g_{0}$. Let $K$ and $g$ denote the true number of communities and the true membership vectors, respectively. Based on the discussions in Hu et al. (2021), the goodness-of-fit test for the SBM is equivalent to testing the following two hypotheses:
  • (I)
    $H_{0k}:K=K_{0}$ v.s. $H_{1k}:K>K_{0}$  and  (II) $H_{0g}:g=g_{0}$ v.s. $H_{1g}:g\not=g_{0}$.
It is worth noting that, for hypothesis test (I), we consider only a one-sided alternative $K>K_{0}$. The main reason is that each SBM with $K$ communities can be split into $K_{0}>K$ communities. Accordingly, we consider a one-sided alternative with nodes being partitioned into less than $K$ communities in hypothesis test (I) to guarantee the power of goodness-of-fit tests (e.g., Lei 2016; Chen and Lei 2018; Wang and Bickel 2017).
To test the null hypotheses in (I) and (II), we employ the maximum likelihood estimation method to estimate the probability matrix $Q=(Q_{g(i)g(j)})$ under the null hypotheses. Consider the probability mass function of the adjacency matrix $A$, $P_{g,Q}(A)=\prod_{1\leq i<j\leq n}Q^{a_{ij}}_{g(i)g(j)}(1-Q^{a_{ij}}_{g(i)g(j)})^{1-a_{ij}}.$ Denote $n_{uv}=\sum_{i=1}^{n}\sum_{j\not=i}\mathbf{1}\{g(i)=u,g(j)=v\}$ and $m_{uv}=\sum_{i=1}^{n}\sum_{j\not=i}a_{ij}\mathbf{1}\{g(i)=u,g(j)=v\}$. Then the corresponding log-likelihood function is $\ell(A|Q,g)=\frac{1}{2}\sum_{u,v=1}^{K}\big{\{}m_{uv}\text{log}~{}Q_{uv}+(n_{uv}-m_{uv})\text{log}(1-Q_{uv})\big{\}}.$ Given the hypothetical number of communities $K_{0}$ and the hypothetical membership vector $g_{0}$, we can obtain the maximum likelihood estimator of $Q$ by maximizing $\ell(A|Q,g_{0})$. It is
$\widehat{Q}_{uv}^{g_{0}}=\left\{\begin{array}[]{l}\frac{\sum_{i\in g_{0}^{-1}(u),j\in g_{0}^{-1}(v)}a_{ij}}{|g_{0}^{-1}(u)||g_{0}^{-1}(v)|},\quad 1\leq u\not=v\leq K_{0},\\ \frac{\sum_{i\not=j\in g_{0}^{-1}(u)}a_{ij}}{|g_{0}^{-1}(u)|(|g_{0}^{-1}(u)|-1)},\quad u=v,\end{array}\right.$ (2.1)
where $g_{0}^{-1}(u)=\{i:1\leq i\leq n,g_{0}(i)=u\}$ and $|g^{-1}_{0}(u)|$ is the number of nodes within the community $u$. Note that, to test the null hypothesis in (I), we replace $g_{0}$ in (2.1) with its consistent estimator $\widehat{g}$; see, e.g., the estimator obtained using Lei and Rinaldo (2015)’s spectral clustering method. We next adopt Hu et al. (2021)’s approach and introduce the maximum entry-wise deviation.
2.2 Maximum entry-wise deviation
To test the null hypotheses in (I) and (II) with a divergent $K$, Hu et al. (2021) proposed the entry-wise deviation,
$\widehat{\rho}_{iv}=\frac{1}{\sqrt{|g_{0}^{-1}(v)/\{i\}|}}\sum_{j\in g_{0}^{-1}(v)/\{i\}}\frac{a_{ij}-\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}}}{\sqrt{\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}}(1-\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}})}},$ (2.2)
where $g_{0}^{-1}(v)/\{i\}$ is the set of nodes contained in $g_{0}^{-1}(v)$ excluding node $i$. They then obtained the maximum entry-wise deviation $L_{n}(K_{0},g_{0})\overset{\Delta}{=}\max_{1\leq i\leq n,1\leq v\leq K_{0}}|\widehat{\rho}_{iv}|$. When the network is dense such that $\min_{i,j}Q_{g(i)g(j)}\geq c$ for a constant $10<c<$ 1, Hu et al. (2021) established the theoretical properties of $L_{n}(K_{0},g_{0})$.
As noted in Section 6 of Hu et al. (2021), their proposed test is not applicable for sparse networks. For the sake of illustration, we consider a sparse network with $c_{\min}n^{-1}\log n\leq\min_{i,j}Q_{g(i)g(j)}\leq\max_{i,j}Q_{g(i)g(j)}\leq c_{\max}n^{-1}\log n$, where $c_{\min}$ and $c_{\max}$ are two finite positive constants. Under this scenario, $\widehat{\rho}_{iv}$ is a summation of $c_{\rho}\log n$ random terms based on (2.2), where $\log n$ is related to the order of network density (i.e., $n^{-1}\log n$), and $c_{\rho}$ is a constant between $c_{\min}$ and $c_{\max}$. Therefore, $\widehat{\rho}_{iv}$ can be expressed as a standard normal distribution $N(0,1)$ plus an additional term $O_{p}(1/\sqrt{\log n})$. Accordingly, it can be demonstrated that the test statistic of Hu et al. (2021), denoted as $L_{n}^{2}(K,g)-2\log(Kn)+\log\log(Kn)$, converges to an extreme value distribution plus an additional term of $O_{p}(1)$. Consequently, the test statistic proposed by Hu et al. (2021) fails to converge to the Gumbel distribution under the sparse network setting, while it is true under the dense network setting demonstrated by Hu et al. (2021). This finding is supported by our simulation experiments given in Subsection 3.2. This motivates us to develop a new test statistic that performs well without regard to the network density (or sparsity).
2.3 Maximum sampling entry-wise deviation test statistic
In this subsection, we propose a novel test statistic that is applicable for both dense and sparse networks. The key idea is to create the maximum entry-wise deviation by integrating the entry-wise deviation $\widehat{\rho}_{iv}$ with the sampling process. Specifically, we calculate the $\widehat{\rho}_{iv}$’s based on Hu et al. (2021). For any fixed community $v\in\{1,\cdots,K_{0}\}$, we randomly draw $B$ samples from $\{\widehat{\rho}_{1v},\cdots,\widehat{\rho}_{nv}\}$ at the $m$-th realization ($m=1,\cdots,M$), and denote them $\{\widehat{\rho}_{m_{1}v},\cdots,\widehat{\rho}_{m_{B}v}\}$. Subsequently, calculate
$\widehat{\gamma}_{mv}=\frac{\widehat{\rho}_{m_{1}v}+\widehat{\rho}_{m_{2}v}+\cdots+\widehat{\rho}_{m_{B}v}}{\sqrt{B}}.$ (2.3)
Repeating this procedure $M$ times for each community $v$, we obtain the sequence of $\widehat{\gamma}_{mv}$’s for $m=1,\cdots,M$. Under the null hypotheses in (I) and (II), one expects all $\widehat{\rho}_{iv}$ values to be small as noted by Hu et al. (2021). Accordingly, $\widehat{\gamma}_{mv}$ should also be small, which motivates us to propose the maximum sampling entry-wise deviation based on the maximum value of $\widehat{\gamma}_{mv}$’s across $m$ and $v$,
$F_{n}(K_{0},g_{0})\overset{\Delta}{=}\max_{1\leq m\leq M,1\leq v\leq K_{0}}|\widehat{\gamma}_{mv}|.$ (2.4)
By setting $B=1$ and choosing $M$ to be sufficiently large such that all the $\widehat{\rho}_{iv}$’s can be selected, $F_{n}(K_{0},g_{0})$ is then identical to $L_{n}(K_{0},g_{0})$.
It is worth noting that the above sampling process can alleviate the negative impact of network sparsity by letting $B$ be divergent in $n$. Considering a sparse network with $c_{\min}n^{-1}\log n\leq\min_{i,j}Q_{g(i)g(j)}\leq\max_{i,j}Q_{g(i)g(j)}\leq c_{\max}n^{-1}\log n$ as discussed in Section 2.2, $\widehat{\rho}_{iv}$ is a summation of $c_{\rho}\log n$ random terms. Accordingly, (2.3) can be expressed as a summation of infinite $c_{\rho}B\log n$ random terms as $B\to\infty$ along with $n$. By the central limit theorem, $\widehat{\gamma}_{mv}$ can be expressed as the standard normal distribution $N(0,1)$ plus a term $O_{p}(1/\sqrt{B\log n})$. By Lemma 1 given in the supplementary material, we can show that $F_{n}(K_{0},g_{0})=O_{p}\big{\{}\sqrt{\log(MK_{0})}\big{\}}\big{\{}O_{p}(1)+O_{p}(1/\sqrt{B\log n})\big{\}}=O_{p}\big{\{}\sqrt{\log(MK_{0})}\big{\}}+o_{p}(1)$ under Conditions (C2) and (C3) below. Thus, the term $O_{p}(1/\sqrt{B\log n})$ is negligible in deriving the asymptotic distribution of $F_{n}(K_{0},g_{0})$. As a result, we expect that $F_{n}(K_{0},g_{0})$ follows an extreme value distribution under certain conditions, and it is applicable to both dense and sparse networks. To rigorously derive the theoretical properties of $F_{n}(K_{0},g_{0})$, we introduce the following technical conditions.
  • (C1)
    Assume that $Q$ has no identical rows. In addition, assume $Q_{g(i)g(j)}\in(0,1)$ for $1\leq i,j\leq n$ uniformly, and $n\min_{i,j}Q_{g(i)g(j)}\to\infty$ as $n\rightarrow\infty$.
  • (C2)
    Assume that there exist two finite positive constants $\xi_{1}$ and $\xi_{2}$ such that
    $\xi_{1}n/K<\min_{1\leq u\leq K}|g^{-1}(u)|\leq\max_{1\leq u\leq K}|g^{-1}(u)|<\xi_{2}n^{2}/(K^{2}\log^{2}n),$
    uniformly for any $n$, where $|g^{-1}(u)|$ is the number of nodes contained in the $u$-th community defined in (2.1).
  • (C3)
    Assume that (i) $B=\max[c_{1}q^{-1/2},c_{2}(n/\log n)^{\xi_{3}-\xi_{c}})]$ for some finite positive constants $c_{1}$, $c_{2}$ and $\xi_{c}$, (ii) $K=o\{(n/\log n)^{\xi_{3}}\}$, and (iii) $MK=O(n^{\xi_{4}})$, where $q=\max_{i,j}Q_{g(i)g(j)}$, and $20<\xi_{c}\leq\xi_{3}<1/$ 2 and $10<\xi_{4}<$ 1 are finite constants. In addition, assume that $\min\{n,M,B\}\rightarrow\infty$.
Condition (C1) requires that the probability matrix $Q$ is identifiable with entries all bounded in $(0,1)$. This condition is similar to, but weaker than, that of Lei (2016), Wang and Bickel (2017), and Hu et al. (2021). Note that if all $Q_{g(i)g(j)}$s have the same order, then $Q_{g(i)g(j)}$ shares the same order with the network density $D_{n}$. Accordingly, this condition immediately implies that the network density $D_{n}$ satisfies $nD_{n}\rightarrow\infty$ and allows the network density to approach zero more slowly than $n^{-1}$ as $n\to\infty$. In addition, if $\widehat{Q}^{g}_{g(i)g(j)}=0$, then the term $\big{(}a_{ij}-\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}}\big{)}/\big{(}\sqrt{\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}}(1-\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}})}\big{)}$ in the summation of (2.2) is removed, and our analysis remains valid. For the sake of simplicity, we omit the case which $Q_{g(i)g(j)}=0$ in this paper. Condition (C2) is a mild condition that requires the size of the smallest community to be at least proportional to $n/K$. Furthermore, Condition (C2) imposes an upper bound on the largest community size, which was also considered by Zhang and Zhou (2016), Gao et al. (2018), and Hu et al. (2021). This condition is used to control the estimation bias between $\widehat{Q}^{g}_{g(i)g(j)}$ and its population version $Q_{g(i)g(j)}$, $\max_{1\leq i\leq n,1\leq v\leq K}\left|\frac{1}{\sqrt{|g^{-1}(v)/\{i\}|}}\sum_{j\in g^{-1}(v)/\{i\}}\frac{Q_{g(i)g(j)}-\widehat{Q}^{g}_{g(i)g(j)}}{\sqrt{Q_{g(i)g(j)}(1-Q_{g(i)g(j)})}}\right|.$ Specifically, the condition ensures that this estimation bias converges to zero in probability. Condition (C3) identifies the orders of $K$, $B$, and $M$. Intuitively, $B$ and $M$ cannot be too large; otherwise, the correlation between $\widehat{\gamma}_{mv}$’s may be large. Note that Condition (C3)-(ii) allows $K$ to grow more slowly along $n$ than the assumption that $K=o{(n/\log^{2}n)}$, which was proposed by Hu et al. (2021). In addition, Han et al. (2023) required $K$ to be a fixed value, although they introduced less stringent density conditions than those in Lei (2016), Wang and Bickel (2017), and Hu et al. (2021). Consequently, in comparison to Lei (2016), Wang and Bickel (2017), Hu et al. (2021), and Han et al. (2023), our method not only exhibits the adaptability to a broad spectrum of network densities, but also allows $K$ to be divergent. Based on the above three conditions, we next obtain the theoretical properties of $T_{n}=F_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0}).$
Theorem 1.
Assume Conditions (C1)-(C3) hold. Under the null hypotheses of $K=K_{0}$ and $g=g_{0}$, we have that, for any $x\in\mathbb{R}$,
$\lim_{n\to\infty}P\Big{\{}T_{n}\leq x\Big{\}}=\exp\big{(}-\frac{1}{\sqrt{\pi}}e^{-x/2}\big{)},$ (2.5)
as $n\to\infty$, where the right hand side of (2.5) is the CDF of the Type-I extreme value distribution (i.e., Gumbel distribution) with $\mu=-2\log(\sqrt{\pi})$ and $\beta=2$.
Applying Theorem 1, we can test the null hypotheses in (I) and (II). To test the null hypothesis in (I), we replace $g_{0}$ with its consistent estimator $\widehat{g}$. This estimator can be obtained regardless the network being dense, $n\min_{i,j}Q_{g(i)g(j)}/\log n\rightarrow\infty$, (see, e.g., Daudin et al. 2008; Lei and Rinaldo 2015; Lei and Zhu 2017; Gao et al. 2017 and Gao et al. 2018) or sparse, $n\max_{i,j}Q_{g(i)g(j)}=O(\log n)$, (see, e.g., Zhang and Zhou 2016; Jing et al. 2022). The resulting test statistic is $T_{n,1}=F_{n}^{2}(K_{0},\widehat{g})-2\log(MK_{0})+\log\log(MK_{0}).$ We can demonstrate theoretically that $T_{n,1}$ has the same asymptotic null distribution as $T_{n}$ under proper conditions; see Corollary S.1 and its proof in Section S.7 of the supplementary material. Consequently, we reject the null hypothesis, $H_{0k}:K=K_{0}$, if $T_{n,1}>q_{\alpha}$, where $q_{\alpha}=-2\log\{-\sqrt{\pi}\log(1-\alpha)\}$ is the $(1-\alpha)$-th quantile of the Type-I extreme distribution with the distribution function $\exp\big{(}-e^{-x/2}/\sqrt{\pi}\big{)}$.
As for testing the null hypothesis in (II), we can directly calculate the test statistic: $T_{n,2}=F_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0}).$ This is because $g_{0}$ can yield its corresponding $K_{0}$. Thus, we can reject the null hypothesis $H_{0g}:g=g_{0}$ when $T_{n,2}>q_{\alpha}$ at the given significance level $\alpha$.
For ease of understanding and practical applications, we present Algorithm 1 below with five steps to calculate the test statistic $T_{n,2}$ for testing (II). To obtain $T_{n,1}$ for testing (I), one can replace $g_{0}$ in $T_{n,2}$ with its consistent estimator $\widehat{g}$ calculated under $K=K_{0}$.
Input: Adjacency matrix $A$, the mapping vector $g_{0}$, the number of communities $K_{0}$ produced by its corresponding $g_{0}$, the number of realizations $M$, the number of random samples $B$, and the significance level $\alpha$.
Output: Test statistic $T_{n,2}$ and the decision in regard to the null hypothesis $H_{0g}$ being rejected or not.
1
2Calculate the probability matrix $\widehat{Q}$ using (2.1).
3Calculate $\widehat{\rho}_{iv}$ using (2.2) for each $1\leq i\leq n$ and $1\leq v\leq K_{0}$.
4At the $m$-th realization for $m=1,\cdots M$, randomly draw $B$ samples from $\{\widehat{\rho}_{1v},\cdots,\widehat{\rho}_{nv}\}$ obtained in Step 2 for each $1\leq v\leq K_{0}$, and calculate $\widehat{\gamma}_{mv}$ using (2.3).
5Find the maximum value of $\widehat{\gamma}_{mv}$ for $1\leq m\leq M$ and $1\leq v\leq K_{0}$, and obtain $F_{n}(K_{0},g_{0})$ using (2.4).
Calculate $T_{n,2}=F_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0})$ and reject the null hypothesis $H_{0g}:g=g_{0}$ if $T_{n,2}>q_{\alpha}$.
Algorithm 1 Maximum Sampling Entry-Wise Deviation Test Statistic
Remark 1.
Although the network considered in this paper requires $n\min_{i,j}Q_{g(i)g(j)}\to\infty$, it can be shown that Theorem 1 is also applicable even under the condition of extreme network sparsity, $n\max_{i,j}Q_{g(i)g(j)}<\infty$, as long as a consistent estimator $\widehat{g}$ can be obtained. For example, a consistent estimator $\widehat{g}$ can be obtained for extremely sparse networks when $K=2$ (see, e.g., Mossel et al. 2015; Mossel et al. 2018; Chin et al. 2015).
Remark 2.
In addition to the testing methods mentioned above, there are selection methods that can be used to choose $K$. Wang and Bickel (2017) proposed a likelihood-based approach to select the optimal SBM, while Li et al. (2020) developed a model-free edge cross-validation procedure for selecting $K$ under SBM. However, these two methods require the stringent conditions: (i) the number of communities $K$ is finite, and (ii) the network is dense with $n\min_{i,j}Q_{g(i)g(j)}/\log n\rightarrow\infty$. To relax condition (ii), Chen and Lei (2018) proposed a network cross-validation (NCV) approach to select $K$, but their method still requires condition (i). In contrast, Chatterjee (2015) introduced an universal singular value thresholding (USVT) estimation method that allows $K$ to diverge. However, his method is not suitable for sparse networks. It is noteworthy that selection- and estimation-type methods are not applicable for testing the model adequacy. Thus, we do not compare our testing method with the aforementioned four methods in simulation and real data analyses.
2.4 The asymptotic power
This subsection investigates the asymptotic power of the proposed test statistic under a class of alternatives modified from Hu et al. (2021). For a given hypothetical membership vector $g_{0}$, we define the probability matrix $Q^{g_{0}}$ as
$Q_{uv}^{g_{0}}=\left\{\begin{array}[]{l}\frac{\sum_{i\in g_{0}^{-1}(u),j\in g_{0}^{-1}(v)}Q_{g(i)g(j)}}{|g_{0}^{-1}(u)||g_{0}^{-1}(v)|},\quad 1\leq u\not=v\leq K_{0},\\ \frac{\sum_{i\not=j\in g_{0}^{-1}(u)}Q_{g(i)g(j)}}{|g_{0}^{-1}(u)|(|g_{0}^{-1}(u)|-1)},\quad u=v.\end{array}\right.$ (2.6)
The differences between the null (i.e., $K=K_{0}$ and $g=g_{0}$) and alternative (i.e., $K>K_{0}$ and $g\not=g_{0}$) hypotheses can be characterized by the differences between $Q=Q^{g}$ and $Q^{g_{0}}$. To derive the asymptotic power under the alternative hypotheses, Hu et al. (2021) introduced an alternative subset defined as $\{(K_{0},g_{0}):K_{0}<K,\ell^{*}(K_{0},g_{0})/\sqrt{\log n}\to\infty\}$, where $\ell^{*}(K_{0},g_{0})=\max_{1\leq i\leq n,1\leq v\leq K_{0}}\Big{|}\sum_{j\in g_{0}^{-1}(v)}\big{(}Q_{g(i)g(j)}-Q^{g_{0}}_{g_{0}(i)g_{0}(j)}\big{)}/\sqrt{|g_{0}^{-1}(v)|}\Big{|}$.
However, $\ell^{*}(K_{0},g_{0})$ is unable to capture the differences between $Q$ and $Q^{g_{0}}$ in sparse networks since both $Q$ and $Q^{g_{0}}$ can decay to zero. For the sake of illustration, we consider sparse networks with $\max_{i,j}Q_{g(i)g_{(}j)}=O(\log n/n)$. Under Condition (C2*) given below, it can be shown that $\ell^{*}(K_{0},g_{0})=O\big{\{}(\max_{u}|g_{0}^{-1}(u)|)^{1/2}\log n/n\big{\}}=O(K_{0}^{-1})$, which implies that $\ell^{*}(K_{0},g_{0})/\sqrt{\log n}\not\to\infty$.
In order to obtain an alternative set that can accommodate sparse networks, we introduce a new measure to gauge the disparity between $Q$ and $Q^{g_{0}}$. Specifically, for any $i\in g^{-1}(u^{*})$ with $u^{*}=1,\cdots,K$, we define $q_{vu^{*}}=\big{|}\frac{1}{\sqrt{\big{|}g_{0}^{-1}(v)/\{i\}\big{|}}}\sum_{j\in g_{0}^{-1}(v)/\{i\}}\frac{Q_{g(i)g(j)}-Q^{g_{0}}_{g_{0}(i)g_{0}(j)}}{\sqrt{Q^{g_{0}}_{g_{0}(i)g_{0}(j)}(1-Q^{g_{0}}_{g_{0}(i)g_{0}(j)})}}\big{|}$, where $v=1,\cdots,K_{0}$. In addition, set
$q_{v}\doteq\sum_{u^{*}=1}^{K}\frac{|g^{-1}(u^{*})|}{n}q_{vu^{*}}.$ (2.7)
We then define the alternative set for $K_{0}$ and $g_{0}$ as follows:
$\mathcal{D}(K,g,Q)=\Big{\{}(K_{0},g_{0}):K_{0}<K,\sqrt{B}\max_{1\leq v\leq K_{0}}q_{v}/\sqrt{\log(MK_{0})}\to\infty\Big{\}},$
where $B$ and $M$ are selected in accordance with Condition (C3). The $\mathcal{D}(K,g,Q)$ indicates that the order of the maximum mean difference between $Q$ and $Q^{g_{0}}$ within the $K_{0}$ blocks is greater than $\sqrt{\log(MK_{0})/B}$ under alternative scenarios.
It is noteworthy that Hu et al. (2021) defined an alternative set $\{(K_{0},g_{0}):K_{0}<K,\ell^{*}(K_{0},g_{0})/\sqrt{\log n}\to\infty\}$. By Condition (C3), we have that $\ell^{*}(K_{0},g_{0})/\sqrt{\log n}\to\infty$ implies $\sqrt{B}\max_{v}q_{v}/\sqrt{\log(MK_{0})}\to\infty$. Therefore, in dense networks, $\{(K_{0},g_{0}):K_{0}<K,\ell^{*}(K_{0},g_{0})/\sqrt{\log n}\to\infty\}$ is a subset of our alternative set $\mathcal{D}(K,g,Q)$. Accordingly, our tests are not only robust against sparsity, but also powerful under Hu et al. (2021)’s alternative set.
To demonstrate the theoretical property of asymptotic power under alternatives in the set $\mathcal{D}(K,g,Q)$, we introduce the following condition on $K_{0}$ and $g_{0}$.
  • (C2*)
    There exist finite constants $\xi_{1}^{*}>0$ and $\xi_{2}^{*}>0$ such that
    $\xi_{1}^{*}n/K_{0}\leq\min_{1\leq u\leq K_{0}}|g_{0}^{-1}(u)|\leq\max_{1\leq u\leq K_{0}}|g_{0}^{-1}(u)|\leq\xi_{2}^{*}n^{2}/(K_{0}^{2}\log^{2}n).$
This condition is mild and analogous to Condition (C2). Basically, it imposes bounds on community sizes for the hypothetical number of communities $K_{0}$ and the hypothetical membership vector $g_{0}$. Consider the test statistic $T_{n,2}=F_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0})$ defined in Subsection 2.3. We next demonstrate its asymptotic property under alternatives.
Theorem 2.
Assume Conditions (C1), (C2*) and (C3) hold. We have that, for any alternative $(K_{0},g_{0})\in\mathcal{D}(K,g,Q)$ and for any finite positive constant $c_{1}$, as $n\to\infty$,
$P\Big{\{}T_{n,2}\geq c_{1}\log(MK_{0})\Big{\}}\to 1.$
Theorem 2 indicates that $T_{n,2}$ diverges at a rate of $\log(MK_{0})$ under the alternative hypothesis in (II) that $(K_{0},g_{0})\in\mathcal{D}(K,g,Q)$. This result is analogous to Theorem 2 in Hu et al. (2021). The main difference is that our proposed $Q$ and $Q^{g_{0}}$ can decay to zero in sparse networks under the alternative set $\mathcal{D}(K,g,Q)$. Accordingly, the resulting test statistic is applicable for both sparse and dense networks. It is worth noting that Theorem 2 can also be used to calculate the power of $T_{n,1}$ under the alternative hypothesis in (I).
Under the null hypothesis in (II), Theorem 1 shows $T_{n,2}=O_{p}(1)$. In contrast, under alternatives, Theorem 2 indicates that the divergence rate of $T_{n,2}$ is $\log(MK_{0})$. Then, applying Condition (C3)-(iii), $MK=O(n^{\xi_{4}})$ with $10<\xi_{4}<$ 1, the rate of $T_{n,2}$ is $\log(n)$, which is the same as that of Hu et al. (2021). From a theoretical perspective, the test statistic $T_{n,2}$, with divergence rate $\log(n)$, is sufficient to separate the null and alternatives asymptotically. In addition, under Conditions (C3)-(ii), $K=o\{(n/\log n)^{\xi_{3}}\}$ with $20<\xi_{3}<1/$ 2, and (C3)-(iii), $MK=O(n^{\xi_{4}})$ with $10<\xi_{4}<$ 1, we have that $M=O(n^{\xi_{4}})$. As a result, $T_{n,2}$ is asymptotically powerful against the alternative set $\mathcal{D}(K,g,Q)$, even for moderately large $M$.
It is noteworthy that we have provided three sufficient conditions to ensure that the alternative $(K_{0},g_{0})$ is contained in $\mathcal{D}(K,g,Q)$; see Proposition 1 in Section S.4 and its proof in Sections S.5 of the supplementary material. In addition, according to both an AE’s and an anonymous referee’s guidance, we have rigorously established the power function of $T_{n,2}$ under some special alternatives in the set $D(K,g,Q)$, and found that the power of $T_{n,2}$ is expected to increase as the network becomes denser (i.e., $Q$ increases), the network size $n$ grows, or the number of communities $K$ decreases; see Proposition 2 and its related discussion in Section S.4, and the proof in Section S.6 of the supplementary material.
2.5 Bootstrap-corrected test statistic
Note that the speed at which $T_{n,1}$ or $T_{n,2}$ converges to the extreme distribution may be slow. We next consider a finite-sample bootstrap-corrected test statistic by generating bootstrap samples to improve the finite sample performance. This bootstrap correction was proposed by Bickel and Sarkar (2016) and also considered by Lei (2016) and Hu et al. (2021). In the remainder of this article, we use generic notation $T_{n}$ to denote either $T_{n,1}$ or $T_{n,2}$ for testing (I) or (II). Given an adjacency matrix $A$, $K=K_{0}$ and $g=g_{0}$, the bootstrap-corrected test statistic $T_{n,boot}$ can be obtained using Algorithm 2 below to test (II), which consists of four steps. Note that $g_{0}$ can be replaced by its consistent estimator $\widehat{g}$ for testing (I).
Input: Adjacency matrix $A$, the mapping vector $g_{0}$, the number of communities $K_{0}$ produced by its corresponding $g_{0}$, the number of realizations $M$, the number of random samples $B$, the number of bootstraps $J$, and the significance level $\alpha$.
Output: Bootstrap-corrected test statistic $T_{n,boot}$, and the decision in regard to the null hypothesis being rejected or not.
1 Calculate $\widehat{Q}$ and then obtain $T_{n}$ from Algorithm 1 using $A$, $g_{0}$, $K_{0}$, $B$ and $M$.
2For $j=1,\cdots,J$, randomly generate a network $A^{(j)}$ based on the stochastic block model $(\widehat{Q},g_{0})$, and obtain $T^{(j)}_{n}$ from Algorithm 1 using $A^{(j)}$, $(\widehat{Q},g_{0})$, $K_{0}$, $B$ and $M$.
3Estimate the location and scale parameters $\widehat{\mu}$ and $\widehat{\beta}$ of the Type-I extreme distribution using the maximum likelihood estimation method given $T^{(j)}_{n}$s for $j=1,\cdots,J$.
Calculate the bootstrap-corrected test statistic $T_{n,boot}=\mu+\beta(\frac{T_{n}-\widehat{\mu}}{\widehat{\beta}})$, where $\mu=-2\log(\sqrt{\pi})$ and $\beta=2$. Reject the null hypothesis if $T_{n,boot}>q_{\alpha}$.
Algorithm 2 Bootstrap-Corrected Test Statistic
We can demonstrate theoretically that the bootstrap-corrected test statistic $T_{n,boot}$ has the same asymptotic null distribution as $T_{n}$; see Corollary S.2 and its proof in Section S.7 of the supplementary material. To examine the finite sample performance of these two tests, simulation studies are reported in the next section. They indicate that $T_{n,boot}$ performs better than $T_{n}$ in both dense and sparse networks.
2.6 An augmented test
As noted by Hu et al. (2021), the power of their proposed test for hypothesis test (I) may be small when the networks are partitioned equally. To improve the power, they proposed a novel technique of adding a new community to the observed network. Consider a planted partition model from an adjacency matrix $A$ with $n$ nodes and $K$ equal-sized communities. Let $K^{+}_{0}=K_{0}+1$. Inspired by the approach of Hu et al. (2021), we propose an augmented test statistic to test (I). In Algorithm 3 below, we first obtain the test statistic $T^{+}_{n}$ to test (II), which consists of seven steps. Then we replace $g_{0}$ in $T^{+}_{n}$ with its consistent estimator $\widehat{g}$ so that it can be used to test (I).
Input: Adjacency matrix $A$, the mapping vector $g_{0}$, the number of communities $K_{0}$ produced by its corresponding $g_{0}$, the number of realizations $M$, the number of random samples $B$, and the significance level $\alpha$.
Output: Augmented test statistic $T^{+}_{n}$ and the decision in regard to the null hypothesis being rejected or not.
1
2Calculate the probability matrix $\widehat{Q}$ using (2.1).
3Add a $K^{+}_{0}$-th community of size $n_{K_{0}^{+}}=\min_{1\leq u\leq K_{0}}|g^{-1}_{0}(u)|/2$ to the observed network with $Q_{K_{0}^{+}K_{0}^{+}}=\max_{1\leq u\leq K_{0}}\widehat{Q}^{g_{0}}_{uu}$ and $Q_{uK_{0}^{+}}=\min_{u\not=v}\widehat{Q}^{g_{0}}_{uv}/2$ for $1\leq u,v\leq K_{0}$.
4Calculate the size $n^{+}=n+n_{K_{0}^{+}}$ and the adjacency matrix $A^{+}$ induced by the $n^{+}$ nodes from Step 2, and then calculate the probability matrix $\widehat{Q}^{+}$ with the membership vector $g^{+}_{0}=(g_{0},\underbrace{K^{+}_{0},\cdots,K^{+}_{0}}_{n_{K^{+}_{0}}})$.
5Calculate $\widehat{\rho}^{+}_{iv}$ using (2.2) with $A^{+}$, $g^{+}_{0}$ and $\widehat{Q}^{+}$, for each $1\leq i\leq n^{+}$ and $1\leq v\leq K^{+}_{0}$.
6At the $m$-th realization for $m=1,\cdots M$, randomly draw $B$ samples from $\{\widehat{\rho}^{+}_{1v},\cdots,\widehat{\rho}^{+}_{n^{+}v}\}$ obtained in Step 4 for each $1\leq v\leq K^{+}_{0}$, and calculate $\widehat{\gamma}^{+}_{mv}$ using (2.3).
7Find the maximum value of $\widehat{\gamma}_{mv}$ for $1\leq m\leq M$ and $1\leq v\leq K^{+}_{0}$ and obtain $F_{n}(K^{+}_{0},g_{0}^{+})$ using (2.4).
Calculate augmented test statistic $T^{+}_{n}=F^{2}_{n}(K^{+}_{0},g_{0}^{+})-2\log(MK^{+}_{0})+2\log\log(MK^{+}_{0})$ and reject the null hypothesis if $T^{+}_{n}>q_{\alpha}$.
Algorithm 3 Augmented Test Statistic
It is worth noting that the above augmented procedure in Step 2 yields a larger disparity measurement, denoted by $q_{vu^{*}}$ in equation (2.7). As a result, $\sqrt{B}\max_{1\leq v\leq K_{0}}q_{v}/\sqrt{\log(MK_{0})}$ becomes larger, which leads to a larger value of $T^{+}_{n}$ so that its power is improved. Additionally, we can demonstrate theoretically that the asymptotic null distribution of $T^{+}_{n}$ is identical to that of $T_{n}$; see Corollary S.3 in Section S.7 of the supplementary material. Moreover, to improve the power of the test statistic by controlling the type-I error close to the 5% nominal level, we can integrate the bootstrap-corrected method introduced in Subsection 2.5 with this augmented test statistic and obtain the augmented bootstrap-corrected statistic $T^{+}_{n,boot}$. Analogously, one can verify that the asymptotic null distribution of $T^{+}_{n,boot}$ is the same as that of $T_{n}$. According to the simulation results, $T^{+}_{n,boot}$ can not only control the size well, but also improve the power. Hence, we include $T^{+}_{n,boot}$ in our simulation studies.
3 Simulation studies
To demonstrate the performance of the proposed test statistic, we conduct simulation studies testing the null hypotheses in (I) and (II) for both dense and sparse networks. To test (I), we employ the simple spectral clustering algorithm to estimate the community membership $\widehat{g}$; see also Lei and Rinaldo (2015). Specifically, for the given adjacency matrix $A$ with $K_{0}$ hypothetical communities, one can recover the community membership $\widehat{g}$ by applying the $k$-means clustering method to the rows of the first $K_{0}$ leading singular vectors of the adjacency matrix $A$. In addition, the network density in our simulation studies is $\tilde{Q}=\sum_{i,j}A_{i,j}/(n\times(n-1))$, where $n$ is the size of network. Then, we set $B=\tilde{Q}^{-1/2}\{n/\log n\}^{1/3}$ and $M=100$. One can easily verify that these definitions of $B$ and $M$ satisfy Condition (C3). The term $\tilde{Q}^{-1/2}$ in $B$ plays a role such that the sparser the network, the more samples are needed in (2.3) to alleviate the sparsity. Our simulation results below indicate that the selected $B$ and $M$ work satisfactorily for both dense and sparse networks.
In studying hypothesis test (I), we not only consider our method, but also include the four competing methods proposed by Hu et al. (2021), Han et al. (2023), Zhang and Amini (2023), and Lei (2016). After including the augmented or the bootstrap-corrected procedure applied to our method, Hu et al. (2021)’s method, and Lei (2016)’s method, we include a total of eight tests in our simulation studies for testing (I). They are $T_{n}$, $T^{+}_{n,boot}$, $T_{Hu}$, $T^{+}_{Hu,boot}$, $T_{Lei}$, $T_{Lei,boot}$, $T_{Han}$, and $T_{Zhang}$, where $T_{Hu}$ and $T^{+}_{Hu,boot}$ are the test statistics proposed by Hu et al. (2021). It is worth noting that the augmented procedure is proposed to test (I), and the methods of Lei (2016), Han et al. (2023), and Zhang and Amini (2023) are not applicable to test (II). Accordingly, we only consider the test statistics $T_{n}$ and $T_{Hu}$ in testing (II). Due to space constraint, all the simulation results for hypothesis test (II) are given in the supplementary material; see Tables S.1-S.2 of Section S.8. In simulation studies, all simulation results are evaluated using 500 realizations.
3.1 Simulations under dense networks
In this subsection, we modify the setting specified by Hu et al. (2021) to construct two dense networks to test hypothesis (I), namely Setting 1.
$\bm{Setting}$ $\bm{1:}$ We set the edge probabilities between any two communities $u$ and $v$ to be $0.1(1+4\times I(u=v))$. As a result, $Q_{g(i)g(j)}=0.1$ if $g(i)\not=g(j)$ and $Q_{g(i)g(j)}=0.5$ otherwise. Thus, $\min_{i,j}Q_{g(i)g(j)}=0.1$, the edge probabilities do not decay with $n$, and the resulting network is a typical dense network. Then, we consider the following two cases: (i) the size of each community $n_{k}=300$ with $K$ communities for $k=1,\cdots,K$, the total number of nodes in this network is $n=Kn_{k}$, and $K_{0}$= 2, 4, 6, 8, and 10; (ii) the size of the network is $n=3,000$ with $K$ blocks, the size of each block is $3000/K$, and $K_{0}$= 3, 5, 10, 15, and 20.
Tables 1-2 present the sizes and powers of all eight tests for cases (i) and (ii), respectively. They show the following findings. First, our proposed tests $T_{n}$ and $T^{+}_{n,boot}$ generally perform better than $T_{Hu}$ and $T^{+}_{Hu,boot}$ in controlling the size, and the sizes of $T_{Hu}$ and $T^{+}_{Hu,boot}$ are significantly and slightly distorted, respectively, when $K>10$. This finding is reasonable, because the sampling process can mitigate the instability of the spectral clustering estimate $\widehat{g}$. Recall that our proposed statistic is constructed by sampling the $B$ entry-wise deviations obtained from Hu et al. (2021). Hence, as $B\rightarrow\infty$, the negative impact due to a small number of poorly estimated $\widehat{g}$ can be ignored. Thus, the performance of the proposed statistic can be dominated by a large number of consistently estimated membership vectors $\widehat{g}$. In addition, both $T_{n}$ and $T^{+}_{n,boot}$ outperform $T_{Lei}$, $T_{Lei,boot}$, $T_{Han}$ and $T_{Zhang}$ in controlling the size. Specifically, the latter four methods exhibit significant size distortions when $K\geq 8$, since their empirical sizes are significantly larger than the nominal level of 5%. Consequently, $T^{+}_{n,boot}$ and $T_{n}$ perform the best overall for controlling the size in dense networks.
Table 1: Setting 1 (i): Proportion of rejections at nominal level $\alpha=0.05$ for hypothesis test $H_{0}:K=K_{0}$ v.s. $H_{1}:K>K_{0}$. Each community has 300 nodes and $Q_{uv}=0.1(1+4\times 1(u=v))$.
$T_{n}$ $T^{+}_{n,boot}$
K 2 4 6 8 10 2 4 6 8 10
$K_{0}=2$ 0.035 0.105 0.225 0.425 0.570 0.045 1 1 1 1
$K_{0}=4$ 0.040 0.095 0.140 0.335 * 0.045 1 1 1
$K_{0}=6$ 0.065 0.090 0.170 * * 0.065 0.975 1
$K_{0}=8$ 0.035 0.065 * * * 0.055 0.840
$K_{0}=10$ 0.050 * * * * 0.050
$T_{Hu}$ $T^{+}_{Hu,boot}$
K 2 4 6 8 10 2 4 6 8 10
$K_{0}=2$ 0.025 0.100 0.385 0.640 0.840 0.055 0.995 1 1 1
$K_{0}=4$ 0.040 0.195 0.410 0.675 * 0.045 1 1 1
$K_{0}=6$ 0.055 0.185 0.385 * * 0.075 1 1
$K_{0}=8$ 0.065 0.160 * * * 0.055 0.995
$K_{0}=10$ 0.070 * * * * 0.065
$T_{Lei}$ $T_{Lei,boot}$
K 2 4 6 8 10 2 4 6 8 10
$K_{0}=2$ 0.050 1 1 1 1 0.045 1 1 1 1
$K_{0}=4$ 0.080 1 1 1 * 0.065 1 1 1
$K_{0}=6$ 0.115 1 1 * * 0.080 1 1
$K_{0}=8$ 0.170 1 * * * 0.135 1
$K_{0}=10$ 0.305 * * * * 0.290
$T_{Zhang}$ $T_{Han}$
K 2 4 6 8 10 2 4 6 8 10
$K_{0}=2$ 0.075 1 1 1 1 0.040 0.715 0.820 0.790 0.815
$K_{0}=4$ 0.040 1 1 1 * 0.065 1 1 0.995
$K_{0}=6$ 0.090 0.995 0.995 * * 0.065 1 1
$K_{0}=8$ 0.100 0.670 * * * 0.095 1
$K_{0}=10$ 0.180 * * * * 0.250
Second, both $T_{n}$ and $T_{Hu}$ exhibit weaker power against alternatives than those of $T_{Lei}$, $T_{Lei,boot}$, $T_{Han}$, and $T_{Zhang}$ in dense networks, and $T_{n}$ is also less powerful than $T_{Hu}$. As expected, the sizes and powers of $T_{n}$ and $T_{Hu}$ can be enhanced by $T^{+}_{n,boot}$ and $T^{+}_{Hu,boot}$, respectively, since they integrate the bootstrap-corrected method with the augmentation procedure. Tables 1-2 demonstrate that the empirical powers of $T^{+}_{n,boot}$ and $T^{+}_{Hu,boot}$ are comparable to those of $T_{Lei}$, $T_{Lei,boot}$, and $T_{Zhang}$, and stronger than that of $T_{Han}$. In addition, $T^{+}_{n,boot}$ is less powerful than $T^{+}_{Hu,boot}$, in particular when $K$ is large, as seen in Table 2. This finding is not surprising since the size of $T^{+}_{Hu,boot}$ is distorted for large $K$. Overall, $T^{+}_{n,boot}$ is a powerful test.
Table 2: Setting 1 (ii): Proportion of rejections at nominal level $\alpha=0.05$ for hypothesis test $H_{0}:K=K_{0}$ v.s. $H_{1}:K>K_{0}$. The network size is $n=3,000$ with equal sized communities, and $Q_{uv}=0.1(1+4\times 1(u=v))$.
$T_{n}$ $T^{+}_{n,boot}$
K 3 5 10 15 20 3 5 10 15 20
$K_{0}=3$ 0.035 0.090 0.470 0.730 0.780 0.030 1 1 1 1
$K_{0}=5$ 0.035 0.145 0.490 0.660 * 0.055 1 1 1
$K_{0}=10$ 0.050 0.120 0.235 * * 0.045 0.645 0.715
$K_{0}=15$ 0.035 0.125 * * * 0.055 0.315
$K_{0}=20$ 0.040 * * * * 0.040
$T_{Hu}$ $T^{+}_{Hu,boot}$
K 3 5 10 15 20 3 5 10 15 20
$K_{0}=3$ 0.020 0.230 0.865 1 1 0.030 1 1 1 1
$K_{0}=5$ 0.025 0.500 0.995 1 * 0.040 1 1 1
$K_{0}=10$ 0.070 0.400 0.980 * * 0.050 0.990 1
$K_{0}=15$ 0.140 0.360 * * * 0.090 0.895
$K_{0}=20$ 0.235 * * * * 0.080
$T_{Lei}$ $T_{Lei,boot}$
K 3 5 10 15 20 3 5 10 15 20
$K_{0}=3$ 0.065 1 1 1 1 0.070 1 1 1 1
$K_{0}=5$ 0.060 1 1 1 * 0.065 1 1 1
$K_{0}=10$ 0.245 1 1 * * 0.235 1 1
$K_{0}=15$ 0.630 1 * * * 0.615 1
$K_{0}=20$ 0.725 * * * * 0.925
$T_{Zhang}$ $T_{Han}$
K 3 5 10 15 20 3 5 10 15 20
$K_{0}=3$ 0.105 1 1 1 1 0.060 1 0.955 0.875 0.750
$K_{0}=5$ 0.070 1 1 1 * 0.070 1 0.975 0.955
$K_{0}=10$ 0.180 0.810 0.915 * * 0.220 1 1
$K_{0}=15$ 0.150 0.830 * * * 0.625 1
$K_{0}=20$ 0.240 * * * * 0.690
In sum, Tables 1-2 demonstrate that $T^{+}_{n,boot}$ performs the best overall as compared with other tests in testing (I), and $T_{n}$ controls size well.
3.2 Simulations under sparse networks
In this subsection, we consider the sparse network setting to test (I), namely Setting 2.
$\bm{Setting}$ $\bm{2:}$ We construct two types of sparse networks as follows: (i) the edge probabilities between any two communities $u$ and $v$ are $2\log\log n(1+4\times I(u=v))/n$. The size of each block is $n_{k}=300$, for $k=1,\cdots,K$, the total number of nodes in the network is $n=Kn_{k}$, and $K_{0}$= 2, 4, 6, 8, and 10; (ii) the edge probabilities between any two communities $u$ and $v$ are $3\log\log n(1+4\times I(u=v))/n$, and the size of the network is $n=3,000$ with $K$ blocks. The network size of each block is $3000/K$, and $K_{0}$= 2, 3, 4, 5, 6, 8, and 10. The above two settings lead to $\max_{i,j}Q_{g(i)g(j)}=10\log\log n/n$ and $\max_{i,j}Q_{g(i)g(j)}=15\log\log n/n$ for networks (i) and (ii), respectively. Accordingly, $\max_{i,j}Q_{g(i)g(j)}=O(n^{-1}\log\log n)$, and it typically yields sparse networks.
Based on the network density conditions imposed by Lei (2016), Hu et al. (2021), Zhang and Amini (2023), and Han et al. (2023), their resulting test statistics, $T_{Lei}$, $T_{Lei,boot}$, $T_{Hu}$, $T^{+}_{Hu,boot}$, $T_{Zhang}$, and $T_{Han}$, are not suitable for networks with sparse density on the order of $O(n^{-1}\log\log n)$. In addition, the methods proposed by Zhang and Amini (2023) and Han et al. (2023) required a fixed value of $K$. Hence, their test statistics may not be applicable to data with a large $K$. Our simulation results in Tables 3–4 corroborate these assertions. Specifically, the empirical sizes of $T_{Lei}$, $T_{Hu}$, and $T^{+}_{Hu,boot}$ are all significantly greater than 5% and suffer from size distortions. Additionally, the empirical sizes of $T_{Zhang}$ are somewhat erratic and none of them approximate to the 5% significance level. As for the tests $T_{Han}$ and $T_{Lei,boot}$, they perform poorly for $K\geq 3$, exhibiting size distortions. On the other hand, the sizes of $T_{n}$ and $T^{+}_{n,boot}$ are close to the nominal level of 5%. Since the sizes of the six competing tests are distorted, it is not sensible to make power comparisons. In sum, under sparse networks, Tables 3-4 indicate that our proposed tests, $T_{n}$ and $T^{+}_{n,boot}$, perform well in comparison with the other six tests, and $T^{+}_{n,boot}$ is more powerful than $T_{n}$, which supports our theoretical results.
Based on Tables 1-4, we conclude that our proposed test $T^{+}_{n,boot}$ performs the best overall in testing (I) under both dense and sparse networks, and $T_{n}$ can control empirical sizes well. In addition, Hu et al. (2021)’s test, $T^{+}_{Hu,boot}$, can be used for dense networks when $K$ is not large.
Table 3: Setting 2 (i): Proportion of rejections at nominal level $\alpha=0.05$ for hypothesis test $H_{0}:K=K_{0}$ v.s. $H_{1}:K>K_{0}$. Each community has 300 nodes and $Q_{uv}=2\log\log n(1+4\times 1(u=v))/n$.
$T_{n}$ $T^{+}_{n,boot}$
K 2 4 6 8 10 2 4 6 8 10
$K_{0}=2$ 0.033 0.075 0.140 0.195 0.245 0.060 0.320 0.515 0.580 0.620
$K_{0}=4$ 0.036 0.085 0.155 0.195 * 0.043 0.380 0.430 0.500
$K_{0}=6$ 0.045 0.090 0.180 * * 0.055 0.235 0.392
$K_{0}=8$ 0.045 0.140 * * * 0.060 0.220
$K_{0}=10$ 0.065 * * * * 0.060
$T_{Hu}$ $T^{+}_{Hu,boot}$
K 2 4 6 8 10 2 4 6 8 10
$K_{0}=2$ 0.610 0.910 0.990 0.990 1 0.970 0.900 0.975 0.985 0.980
$K_{0}=4$ 1 1 1 1 * 0.900 0.990 0.995 0.995
$K_{0}=6$ 1 1 1 * * 0.990 1 1
$K_{0}=8$ 1 1 * * * 0.996 1
$K_{0}=10$ 1 * * * * 1
$T_{Lei}$ $T_{Lei,boot}$
K 2 4 6 8 10 2 4 6 8 10
$K_{0}=2$ 1 1 1 1 1 0.050 1 0.265 0.487 1
$K_{0}=4$ 1 1 1 1 * 0.025 1 0.337 0.644
$K_{0}=6$ 1 1 1 * * 0.226 0.500 0.556
$K_{0}=8$ 1 1 * * * 0.274 0.543
$K_{0}=10$ 1 * * * * 0.221
$T_{Zhang}$ $T_{Han}$
K 2 4 6 8 10 2 4 6 8 10
$K_{0}=2$ 0.080 0.980 0.210 0.165 0.125 0.030 0.745 0.910 0.950 0.315
$K_{0}=4$ 0.125 0.115 0.105 0.160 * 0.635 0.830 0.980 0.955
$K_{0}=6$ 0.020 0.065 0.080 * * 0.960 0.930 0.935
$K_{0}=8$ 0.008 0.070 * * * 0.930 0.875
$K_{0}=10$ 0.003 * * * * 0.590
Table 4: Setting 2 (ii): Proportion of rejections at nominal level $\alpha=0.05$ for hypothesis test $H_{0}:K=K_{0}$ v.s. $H_{1}:K>K_{0}$. The network size is $n=3,000$ with equal sized communities, and $Q_{uv}=3\log\log n(1+4\times 1(u=v))/n$.
$T_{n}$ $T^{+}_{n,boot}$
K 2 3 4 5 6 8 10 2 3 4 5 6 8 10
$K_{0}=2$ 0.043 0.060 0.110 0.170 0.265 0.325 0.360 0.060 0.650 0.685 0.750 0.780 0.800 0.825
$K_{0}=3$ 0.055 0.095 0.115 0.135 0.205 0.320 * 0.050 0.560 0.685 0.700 0.765 0.800
$K_{0}=4$ 0.055 0.065 0.130 0.185 0.210 * * 0.060 0.420 0.485 0.565 0.620
$K_{0}=5$ 0.065 0.110 0.160 0.195 * * * 0.040 0.365 0.455 0.550
$K_{0}=6$ 0.070 0.105 0.160 * * * * 0.035 0.305 0.355
$K_{0}=8$ 0.065 0.110 * * * * * 0.070 0.310
$K_{0}=10$ 0.070 * * * * * * 0.060
$T_{Hu}$ $T^{+}_{Hu,boot}$
K 2 3 4 5 6 8 10 2 3 4 5 6 8 10
$K_{0}=2$ 0.537 0.960 0.970 0.985 0.995 1 1 0.993 0.930 0.950 0.970 0.995 1 1
$K_{0}=3$ 0.975 1 1 1 1 1 * 0.870 0.980 0.980 0.995 1 1
$K_{0}=4$ 1 1 1 1 1 * * 0.815 0.975 1 1 1
$K_{0}=5$ 1 1 1 1 * * * 0.955 1 1 1
$K_{0}=6$ 1 1 1 * * * * 1 1 1
$K_{0}=8$ 1 1 * * * * * 1 1
$K_{0}=10$ 1 * * * * * * 1
$T_{Lei}$ $T_{Lei,boot}$
K 2 3 4 5 6 8 10 2 3 4 5 6 8 10
$K_{0}=2$ 1 1 1 1 1 1 1 0.050 1 1 1 0.495 0.245 0.517
$K_{0}=3$ 1 1 1 1 1 1 * 0.100 1 1 0.140 0.368 0.614
$K_{0}=4$ 1 1 1 1 1 * * 0.080 1 0.250 0.166 0.399
$K_{0}=5$ 1 1 1 1 * * * 0.110 1 0.550 0.156
$K_{0}=6$ 1 1 1 * * * * 0.165 0.518 0.477
$K_{0}=8$ 1 1 * * * * * 0.388 0.473
$K_{0}=10$ 1 * * * * * * 0.343
$T_{Zhang}$ $T_{Han}$
K 2 3 4 5 6 8 10 2 3 4 5 6 8 10
$K_{0}=2$ 0.130 1 1 1 0.995 0.355 0.145 0.025 0.795 0.760 0.620 0.620 0.525 0.632
$K_{0}=3$ 0.075 1 1 0.995 0.267 0.313 * 0.085 0.605 0.805 0.770 0.848 0.900
$K_{0}=4$ 0.025 1 0.975 0.133 0.153 * * 0.150 0.745 0.817 0.945 0.985
$K_{0}=5$ 0.010 0.420 0.193 0.113 * * * 0.360 0.910 0.940 0.960
$K_{0}=6$ 0.010 0.080 0.080 * * * * 0.865 0.965 0.965
$K_{0}=8$ 0.030 0.070 * * * * * 0.915 0.930
$K_{0}=10$ 0.020 * * * * * * 0.890
Remark 3. We also construct simulation studies under Settings 1 and 2 to test hypothesis (II). Due to space constraint, all of the simulation results for testing hypothesis (II) are given in Tables S.1-S.2 of Section S.8 in the supplementary material. The simulation results in Tables S.1-S.2 are consistent with those findings in Tables 1-4. Under Setting 1, Table S.1 indicates that our proposed statistic $T_{n}$ is superior to $T_{Hu}$ in the size control, and both $T_{n}$ and $T_{Hu}$ are powerful against the alternatives with the empirical powers all equal to 1. Under Setting 2, Table S.2 shows that the empirical sizes of $T_{n}$ are close to the nominal level of 5% and the empirical powers of $T_{n}$ are all 1, while $T_{Hu}$ fails to control size well. In sum, when testing hypothesis (II), $T_{n}$ performs well and is superior to $T_{Hu}$ in the size control under both dense and sparse networks.
Remark 4. It is worth noting that the condition $D_{n}=O(n^{-1}\log n)$ has been considered in obtaining theoretical properties of various tests; see, e.g., the $T_{Zhang}$ test. According to this condition, we construct additional simulation studies under Setting 2 to test hypotheses (I) and (II), except that the edge probabilities between any two communities $u$ and $v$ are adjusted to $2\log n(1+4\times I(u=v))/n$ and $3\log n(1+4\times I(u=v))/n$, respectively, for cases (i) and (ii). Additional simulation results under this setting are presented in Tables S.3-S.5 of Section S.9 in the supplementary material. The simulation results in Tables S.3-S.5 are consistent with those findings in Tables 3-4. In sum, when testing hypothesis (I), our proposed test statistics $T_{n}$ and $T^{+}_{n,boot}$ outperform all competing methods, and $T^{+}_{n,boot}$ is more powerful than $T_{n}$ and performs the best overall. When testing hypothesis (II), $T_{n}$ is superior to $T_{Hu}$.
4 Real data analyses
To demonstrate the usefulness of our proposed method for both dense and sparse networks, we analyze two real datasets. Based on simulation studies, we find that our proposed method and Hu et al. (2021)’s method generally outperform those methods proposed by Lei (2016), Han et al. (2023), and Zhang and Amini (2023). Thus, in our real data analyses, we only consider our proposed method and Hu et al. (2021)’s method. The first dataset is the international trade data, which has been studied by Westveld and Hoff (2011), Lei (2016), and Hu et al. (2021). This data set contains yearly trade information of 58 countries. Hu et al. (2021) constructed a network with density of 48.1% (=1590/(58×57)×1590/(58$\times$ 57$\times$1590 / ( 58 × 57 ) × 100%), and it is a typical dense network. As noticed by Hu et al. (2021), the augmentation procedure may not work well for this data set because the size of the network is small and the added community may have less than or equal to two nodes. Therefore, we apply the test statistics, $T_{n,boot}$ and $T_{Hu,boot}$, to analyze the international trade network. The second data set is the co-authorship network dataset collected by Ji and Jin (2016). This network has 3,607 nodes and the density of this network is approximately equal to 0.012% (=1561/(3607×3606)×1561/(3607$\times$ 3606$\times$1561 / ( 3607 × 3606 ) ×100%), which makes it a sparse network. The simulation results indicate that $T^{+}_{n,boot}$ and $T^{+}_{Hu,boot}$ perform better than $T_{n}$ and $T_{Hu}$, respectively. Hence, we employ them to study the co-authorship network. However, the augmented tests are not applicable for testing $H_{0}:K_{0}=1$ since it is infeasible to specify the connection probability between communities. Thus, $T_{n,boot}$ and $T_{Hu,boot}$ are considered in this case. Finally, the number of bootstraps used in both examples is $M=100$, and the number of samples drawn from $n$ nodes is $B=\tilde{Q}^{-1/2}\{n/\log n\}^{1/3}$ as given in simulation studies.
Based on the simulation results presented in Tables 1-4, our proposed method not only is superior to or comparable with the methods proposed by Lei (2016), Hu et al. (2021), Han et al. (2023) and Zhang and Amini (2023) in dense networks, but also outperforms them in sparse networks. Hence, our method is robust against network density (or sparsity). In empirical analyses, we consider two examples corresponding to dense and sparse networks. Based on theoretical and simulation results, our proposed methods are applicable for the two data sets.
4.1 Dense network case: International trade data
This dataset contains yearly trade information for $n=58$ countries from 1981 to 2000. For any given year, $\text{Trade}_{ij}$ represents the value of exports from country $i$ to country $j$; further descriptions can be found in Westveld and Hoff (2011). Following Saldana et al. (2017) and Hu et al. (2021), we consider the international trade data in 1995, and construct the network from the 58 countries based on $\text{Trade}_{ij}$ as follows. Each node represents a country, and define $W_{ij}=\text{Trade}_{ij}+\text{Trade}_{ji}$. Then, the adjacency matrix $A=(a_{ij})$ is defined such that $a_{ij}=1$ if $W_{ij}>W_{0.5}$ and $a_{ij}=0$ otherwise, where $W_{0.5}$ is the 50% quantile of the $W_{ij}$ values. As discussed in Saldana et al. (2017), it is possible to decompose the network into 3, 7, and 10 communities by three different methods. However, for different choice of $K$, the resulting communities from these 58 countries vary. Therefore, determining the appropriate number of communities is critical to label communities and make meaningful interpretations for 58 countries. To this end, we employ the statistics $T_{n,boot}$ and $T_{Hu,boot}$ to test the three hypotheses $H_{0}:K=3$, $H_{0}:K=7$, and $H_{0}:K=10$, respectively.
Given the 5% significance level, we apply our Theorem 1 and Hu et al. (2021)’s Theorem 1 to obtain the critical values of $T_{n,boot}$ and $T_{Hu,boot}$, which are $t_{n,0.95}=4.792$ and $t_{Hu,0.95}=3.413$, respectively. For $K_{0}=3,7,$ and 10, Hu et al. (2021) calculated their corresponding test statistics, $T_{Hu,boot}=44.28,3.33$ and 13.44. Comparing them with the critical value, 3.413, they concluded that $K=7$ is a reasonable choice. We next compute our proposed test statistics, $T_{n,boot}=15.02,3.07,$ and 16.76, for $K_{0}=3,7,$ and 10, respectively. Comparing them with the critical value, 4.792, we find that $T_{n,boot}$ does not reject the hypothesis $H_{0}:K=7$, whereas it rejects the hypotheses $H_{0}:K=3$ and $H_{0}:K=10$. Hence, $T_{n,boot}$ yields the same finding as that of Hu et al. (2021).
Based on our test results, the 58 countries can be decomposed into seven blocks. Accordingly, we apply the spectral clustering algorithm of Lei and Rinaldo (2015) and classify 58 countries into seven communities. The results are depicted in Figure S.1 of the supplementary material, which allows us to make the following interpretations regarding the seven communities. The first community marked in purple comprises 4 countries including Cyprus, Egypt and Greece, located in the Mediterranean. The second community marked in blue-purple comprises 6 countries including Argentina, Bolivia and Chile, located in South America. The third community marked in green includes 3 countries, Algeria, Morocco and Tunisia, all located in North Africa. The fourth community marked in blue-green consists of 12 countries including Italy, Spain and Portugal, and most of these countries are located in Europe. It seems that these four communities are identified by regional location, as these countries may be focusing in part on a local trading bloc that provides preferential access. The fifth community marked in blue comprises 9 countries including Norway, Singapore and Denmark, most of which have a high average GDP. The sixth community marked in yellow consists of 13 countries, including USA, Germany, and Japan, most of which are highly influential over global trading volume. The seventh community marked in orange comprises 11 countries including Australia, Canada and the United Kingdom, many of which serve as important hubs of global trade and have GDPs exceeding the global average. The last three communities seem to be identified by these countries’ GDP or trading volume; hence, these countries play a pivotal role in global commerce. This decomposition can provide a better understanding of supply and demand via international trade.
4.2 Sparse network case: Co-authorship data
We next analyze the co-authorship dataset collected by Ji and Jin (2016). This dataset contains all research papers published from 2003 to the first half of 2012 in four of the top statistical journals: Annals of Statistics (AoS), Biometrika, the Journal of American Statistical Association (JASA), and the Journal of Royal Statistical Society (Series B) (JRSS-B). The dataset contains $n=3,607$ authors. According to Ji and Jin (2016)’s Coauthorship Network (A), we construct the network as follows. Each node represents a author. For any author $i$ and author $j$ ($i\not=j$), we define $a_{ij}=1$ for $1\leq i,j\leq n$ if and only if they have co-authored at least two papers during the study period.
The density of this network is approximately 0.012%, which is smaller than 0.23% (i.e., $\log(n)/n$ with $n=3,607$). As mentioned at the beginning of Section 4, this network is considered sparse. Utilizing this network data, Ji and Jin (2016) employed various community detection methods to identify meaningful communities or research groups. Specifically, based on their result in Table 1, the dataset is comprised of eight communities. In their studies, the number of communities corresponds to the number of major research topics or research institutions. Therefore, identifying the number of communities is crucial for their analyses.
Using this sparse network data, we test the hypothetical number of communities $K_{0}\in[1,8]$. Given the 5% significance level, we have $t_{n,0.95}=4.792$ and $t_{Hu,0.95}=3.413$ as given in the above example. In testing $H_{0}:K=1$, both $T_{n,boot}$ and $T_{Hu,boot}$ are larger than 400, which results in rejecting the null hypothesis. For $K_{0}\in[2,8]$, we have $T^{+}_{Hu,boot}>7$, which results in rejecting all of their corresponding hypotheses. This finding is not surprising since $T^{+}_{Hu,boot}$ fails to converge to an extreme-value distribution for sparse networks. In testing $K_{0}\in[2,8]$, we calculate $T^{+}_{n,boot}$=14.81, 15.80, 4.23, -1.53, 3.75, -2.52, and 3.08. As a result, $T^{+}_{n,boot}$ does not reject each of those hypotheses from $H_{0}:K=4$ to $H_{0}:K=8$ at the 5% significance level. Hence, $T^{+}_{n,boot}$ can identify $K=4$ to $K=8$ communities for all authors.
Based on Ji and Jin (2016)’s finding in their Table 1, $K=8$ is sensible. As for $K=4$, we can identify the following four communities in the Coauthorship Network (A): “High-Dimensional Data”, “Machine Learning”, “Experimental Design” and “Prestige and Non-Ivy League Universities”. Please note that the community of “High-Dimensional Data” and “Machine Learning” are merged from “High-Dimensional Data Analysis and Dimension Reduction” and “Theoretical Machine Learning and Quantile Regression”, respectively, in Table 1 of Ji and Jin (2016). When $K=5,6,$ and $7$, we can analogously obtain their corresponding communities by combining similar features from $K=8$. As discussed in Ji and Jin (2016), such decompositions can not only identify the hot areas and primary authors in research communities, but also enhance understanding of the research habits and trends of scientific researchers. Consequently, our proposed test is applicable for detecting community structures in this sparse network data, which provides more sensible options for practitioners to select the number of communities.
In sum, our proposed test and Hu et al. (2021)’s test lead to the same result in the dense network example, and we also provide more interpretations based on the Editor’s comments. In the sparse network example, our test successfully detects a reasonable number of network communities, whereas Hu et al. (2021)’s test rejects all possible hypotheses and fails to yield any conclusion. Consequently, our proposed method is applicable for both data sets.
5 Concluding remarks
In this paper, we combine Hu et al. (2021)’s approach with a sampling process to develop a novel test statistic to test the SBM for both dense and sparse networks. We demonstrate theoretically that the proposed test statistic converges to the Type-I extreme value distribution under the null hypothesis, regardless of the network structure. In addition, we have established the asymptotic power properties and introduced the bootstrap-corrected method and an augmented test statistic. Moreover, simulation studies support our theoretical findings and two empirical examples illustrate the usefulness of the proposed method.
It is of interest to note that our proposed method can be modified to accommodate the degree-corrected stochastic block model (DCSBM; Karrer and Newman 2011). Compared with the SBM, the DCSBM includes a set of node degree parameters $\omega=(\omega_{i})_{1\leq i\leq n}$ for measuring the degree variation of nodes and assumes that $P(A_{ij}=1|Q,g,\omega)=\omega_{i}\omega_{j}Q_{g(i)g(j)}$. The log-likelihood function under the DCSBM is, $\ell(A|Q,g,\omega)=\sum_{i=1}^{n}d_{i}\log\omega_{i}+2^{-1}\sum_{u,v=1}^{K}(m_{uv}\log Q_{uv}-n_{uv}\log Q_{uv})$, where $m_{uv}$ and $n_{uv}$ are defined in Subsection 2.1, and $d_{i}=\sum_{j=1}^{n}A_{ij}$. The maximum likelihood estimators $\widehat{Q}^{g_{0}}$ and $\widehat{\omega}$ can then be obtained by maximizing $\ell(A|Q,g_{0},\omega)$ for the given hypothetical values of $K_{0}$ and $g_{0}$. Accordingly, our proposed method can be applied to the DCSBM by first replacing $\widehat{Q}^{g_{0}}_{g_{0}(i)g_{0}(j)}$ with $\widehat{\omega}_{i}\widehat{\omega}_{j}\widehat{Q}^{g_{0}}_{g_{0}(i)g_{0}(j)}$ in (2.2), and then obtaining statistics analogous to those in (2.3) and (2.4). The asymptotic null distribution and the power of the resulting test statistic under the DCSBM can be demonstrated theoretically. Since the theoretical properties and numerical studies under the DCSBM are almost the same as those of the SBM, we do not present them in the paper.
Rerefences
12013Agarwal et al.Agarwal et al. (2013)agarwal2013traffic Agarwal, S., Kodialam, M., and Lakshman, T. (2013), “Traffic Engineering in Software Defined Networks,” 2013 Proceedings IEEE INFOCOM, 2211–2219. 22013Amini et al.Amini et al. (2013)amini2013pseudo Amini, A. A., Chen, A., Bickel, P. J., and Levina, E. (2013), “Pseudo-Likelihood Methods for Community Detection in Large Sparse Networks,” The Annals of Statistics, 41, 2097–2122. 32009Bickel and ChenBickel and Chen (2009)bickel2009nonparametric Bickel, P. J., and Chen, A. (2009), “A Nonparametric View of Network Models and Newman–Girvan and Other Modularities,” Proceedings of the National Academy of Sciences, 106, 21068–21073. 42016Bickel and SarkarBickel and Sarkar (2016)bickel2016hypothesis Bickel, P. J., and Sarkar, P. (2016), “Hypothesis Testing for Automated Community Detection in Networks,” Journal of the Royal Statistical Society Series B: Statistical Methodology, 78, 253–273. 52015ChatterjeeChatterjee (2015)chatterjee2015matrix Chatterjee, S. (2015), “Matrix Estimation by Universal Singular Value Thresholding,” The Annals of Statistics, 43, 177–214. 62018Chen and LeiChen and Lei (2018)chen2018network Chen, K., and Lei, J. (2018), “Network Cross-Validation for Determining the Number of Communities in Network Data,” Journal of the American Statistical Association, 113, 241–251. 72015Chin et al.Chin et al. (2015)chin2015stochastic Chin, P., Rao, A., and Vu, V. (2015), “Stochastic Block Model and Community Detection in Sparse Graphs: A Spectral Algorithm with Optimal Rate of Recovery,” Conference on Learning Theory, 391–423. 82012Choi et al.Choi et al. (2012)choi2012stochastic Choi, D. S., Wolfe, P. J., and Airoldi, E. M. (2012), “Stochastic Blockmodels with a Growing Number of Classes,” Biometrika, 99, 273–284. 92008Daudin et al.Daudin et al. (2008)daudin2008mixture Daudin, J. J., Picard, F., and Robin, S. (2008), “A Mixture Model for Random Graphs,” Statistics and Computing, 18, 173–183. 102019Di Maggio et al.Di Maggio et al. (2019)di2019relevance Di Maggio, M., Franzoni, F., Kermani, A., and Sommavilla, C. (2019), “The Relevance of Broker Networks for Information Diffusion in the Stock Market,” Journal of Financial Economics, 134, 419–446. 112017Gao et al.Gao et al. (2017)gao2017achieving Gao, C., Ma, Z., Zhang, A. Y., and Zhou, H. H. (2017), “Achieving Optimal Misclassification Proportion in Stochastic Block Models,” The Journal of Machine Learning Research, 18, 1980–2024. 122018Gao et al.Gao et al. (2018)gao2018community Gao, C., Ma, Z., Zhang, A. Y., and Zhou, H. H. (2018), “Community Detection in Degree-Corrected Block Models,” The Annals of Statistics, 46, 2153–2185. 132023Han et al.Han et al. (2023)han2023universal Han, X., Yang, Q., and Fan, Y. (2023), “Universal Rank Inference via Residual Subsampling with Application to Large Networks,” The Annals of Statistics, 51, 1109–1133. 141983Holland et al.Holland et al. (1983)holland1983stochastic Holland, P. W., Laskey, K. B., and Leinhardt, S. (1983), “Stochastic Blockmodels: First Steps,” Social Networks, 5, 109–137. 152021Hu et al.Hu et al. (2021)hu2021using Hu, J., Zhang, J., Qin, H., Yan, T., and Zhu, J. (2021), “Using Maximum Entry-Wise Deviation to Test the Goodness of Fit for Stochastic Block Models,” Journal of the American Statistical Association, 116, 1373–1382. 162016Ji and JinJi and Jin (2016)ji2016coauthorship Ji, P., and Jin, J. (2016), “Coauthorship and Citation Networks for Statisticians,” The Annals of Applied Statistics, 10, 1779–1812. 172015JinJin (2015)jin2015fast Jin, J. (2015), “Fast Community Detection by SCORE,” The Annals of Statistics, 43, 57–89. 182022Jing et al.Jing et al. (2022)jing2022community Jing, B. Y., Li, T., Ying, N., and Yu, X. (2022), “Community Detection in Sparse Networks Using the Symmetrized Laplacian Inverse Matrix (Slim),” Statistica Sinica, 32. 192011Karrer and NewmanKarrer and Newman (2011)karrer2011stochastic Karrer, B., and Newman, M. E. (2011), “Stochastic Blockmodels and Community Structure in Networks,” Physical Review E, 83, 016107. 202016Karwa et al.Karwa et al. (2016)karwa2016exact Karwa, V., Pati, D., Petrovic, S., Solus, L., Alexeev, N., Raic, M., Wilburne, D., Williams, R., and Yan, B. (2016), “Exact Tests for Stochastic Block Models,” arXiv preprint arXiv:1612.06040. 212011Katona et al.Katona et al. (2011)katona2011network Katona, Z., Zubcsek, P. P., and Sarvary, M. (2011), “Network Effects and Personal Influences: The Diffusion of an Online Social Network,” Journal of Marketing Research, 48, 425–443. 222015LawyerLawyer (2015)lawyer2015understanding Lawyer, G. (2015), “Understanding the Influence of All Nodes in a Network,” Scientific Reports, 5, 8665. 232019Lee and WilkinsonLee and Wilkinson (2019)lee2019review Lee, C., and Wilkinson, D. J. (2019), “A Review of Stochastic Block Models and Extensions for Graph Clustering,” Applied Network Science, 4, 1–50. 242016LeiLei (2016)lei2016goodness Lei, J. (2016), “A Goodness-of-Fit Test for Stochastic Block Models,” The Annals of Statistics, 44, 401–424. 252015Lei and RinaldoLei and Rinaldo (2015)lei2015consistency Lei, J., and Rinaldo, A. (2015), “Consistency of Spectral Clustering in Stochastic Block Models,” The Annals of Statistics, 43, 215–237. 262017Lei and ZhuLei and Zhu (2017)lei2017generic Lei, J., and Zhu, L. (2017), “Generic Sample Splitting for Refined Community Recovery in Degree Corrected Stochastic Block Models,” Statistica Sinica, 1639–1659. 272020Li et al.Li et al. (2020)li2020network Li, T., Levina, E., and Zhu, J. (2020), “Network Cross-Validation by Edge Sampling,” Biometrika, 107, 257–276. 282015Mossel et al.Mossel et al. (2015)mossel2015reconstruction Mossel, E., Neeman, J., and Sly, A. (2015), “Reconstruction and Estimation in the Planted Partition Model,” Probability Theory and Related Fields, 162, 431–461. 292018Mossel et al.Mossel et al. (2018)mossel2018proof Mossel, E., Neeman, J., and Sly, A. (2018), “A Proof of the Block Model Threshold Conjecture,” Combinatorica, 38, 665–708. 302018Musumeci et al.Musumeci et al. (2018)musumeci2018overview Musumeci, F., Rottondi, C., Nag, A., Macaluso, I., Zibar, D., Ruffini, M., and Tornatore, M. (2018), “An Overview on Application of Machine Learning Techniques in Optical Networks,” IEEE Communications Surveys &\&& Tutorials, 21, 1383–1408. 312006NewmanNewman (2006)newman2006modularity Newman, M. E. (2006), “Modularity and Community Structure in Networks,” Proceedings of the National Academy of Sciences, 103, 8577–8582. 322001Nowicki and SnijdersNowicki and Snijders (2001)nowicki2001estimation Nowicki, K., and Snijders, T. A. B. (2001), “Estimation and Prediction for Stochastic Blockstructures,” Journal of the American Statistical Association, 96, 1077–1087. 332011Rohe et al.Rohe et al. (2011)rohe2011spectral Rohe, K., Chatterjee, S., and Yu, B. (2011), “Spectral Clustering and the High-Dimensional Stochastic Blockmodel,” The Annals of Statistics, 39, 1878–1915. 342017Saldana et al.Saldana et al. (2017)saldana2017many Saldana, D. F., Yu, Y., and Feng, Y. (2017), “How Many Communities Are There?,” Journal of Computational and Graphical Statistics, 26, 171–181. 352015Sarkar and BickelSarkar and Bickel (2015)sarkar2015role Sarkar, P., and Bickel, P. J. (2015), “Role of Normalization in Spectral Clustering for Stochastic Blockmodels,” The Annals of Statistics, 43, 962–990. 361997Snijders and NowickiSnijders and Nowicki (1997)snijders1997estimation Snijders, T. A., and Nowicki, K. (1997), “Estimation and Prediction for Stochastic Blockmodels for Graphs with Latent Block Structure,” Journal of Classification, 14, 75–100. 372018Tabassum et al.Tabassum et al. (2018)tabassum2018social Tabassum, S., Pereira, F. S., Fernandes, S., and Gama, J. (2018), “Social Network Analysis: An Overview,” Wiley Interdisciplinary Reviews: Data Mining and Knowledge Discovery, 8, e1256. 382010ValenteValente (2010)valente2010social Valente, T. W. (2010), Social Networks and Health: Models, Methods, and Applications, Oxford University Press. 392017Wang and BickelWang and Bickel (2017)wang2017likelihood Wang, Y. R., and Bickel, P. J. (2017), “Likelihood-Based Model Selection for Stochastic Block Models,” The Annals of Statistics, 45, 500–528. 402011Westveld and HoffWestveld and Hoff (2011)westveld2011mixed Westveld, A. H., and Hoff, P. D. (2011), “A Mixed Effects Model for Longitudinal Relational and Network Data, with Applications to International Trade and Conflict,” The Annals of Applied Statistics, 5, 843–872. 412022Wu et al.Wu et al. (2022)wu2022inward Wu, Y., Lan, W., Zou, T., and Tsai, C. L. (2022), “Inward and Outward Network Influence Analysis,” Journal of Business &\&& Economic Statistic, 40, 1617–1628. 422016Zhang and ZhouZhang and Zhou (2016)zhang2016minimax Zhang, A. Y., and Zhou, H. H. (2016), “Minimax Rates of Community Detection in Stochastic Block Models,” The Annals of Statistics, 44, 2522–2280. 432023Zhang and AminiZhang and Amini (2023)zhang2020adjusted Zhang, L., and Amini, A. A. (2023), “Adjusted Chi-Square Test for Degree-Corrected Block Models,” The Annals of Statistics, 51, 2366–2385. 442012Zhao et al.Zhao et al. (2012)zhao2012consistency Zhao, Y., Levina, E., and Zhu, J. (2012), “Consistency of Community Detection in Networks Under Degree-Corrected Stochastic Block Models,” The Annals of Statistics, 40, 2266–2292.